Could you please help me solve system of equations
5x + 3y =5(4-x) AND 2x - 5y = -(y + 11)/2

Answer by checkley71(8403)

5X+3Y=5(4-X)
5X+3Y=20-5X
3Y=-5X-5X+20
3Y=-10X+20
Y=-10X/3+20/3
---------------------------------------
2X-5Y=-(Y+11)/2
2X+5Y=-Y-5.5
5Y+Y=-2X-5.5
6Y=-2X-5.5
Y=-2X/6-5.5/6
Y=-X/3-5.5/6